OpenStudy (anonymous):

John’s salary was increased by 25% in the 1st year and by 20% in the second year by what % should his alary be increased in the third year so that it becomes twice of the iitial value?

OpenStudy (anonymous):

@nbouscal

OpenStudy (anonymous):

Let's do this the same way we did the other one, by using an example so we can think about it easier. Say John's salary was 100 to start with. After the first year, it was 125. Then after the second year, it was 150 (because 20% of 125 is 25, and 125+25=150). So what % of 150 do we need to increase it by to make it twice the initial value, which would be 200?

OpenStudy (anonymous):

400/3?

OpenStudy (anonymous):

How did you get that? We want to get from 150 to 200, so we need to increase it by 50. So what percent of 150 is 50?

OpenStudy (anonymous):

Oh I think I see what you did. 200 is already twice the initial value of 100, you don't need to double it again.

OpenStudy (anonymous):

Wait 100 increased by 25% is 125 then 125 increased by 20% is 150 now 150 wants to ge to 200 so 150 (x/100) = 200

OpenStudy (anonymous):

150*(1+x/100)=200 actually, but yes.

OpenStudy (anonymous):

25%?

OpenStudy (anonymous):

Not quite, 25% of 150 is only 37.5, so it would only get you up to 187.5

OpenStudy (anonymous):

I dont understand

OpenStudy (anonymous):

@nbouscal

OpenStudy (anonymous):

You're trying to get from 150 to 200. So that's a difference of 50. So you need to find out what percentage 50 is of 150.

OpenStudy (anonymous):

Is it 300?

OpenStudy (anonymous):

@hashsam1 can u help?

OpenStudy (anonymous):

300%? 300% of 150 is 450. If it's easier, figure out what fraction you can multiply 150 by to get 50, and then convert the fraction into a percentage.

OpenStudy (anonymous):

Can u simply tell me what it is? |I'm absolutely clueless

OpenStudy (anonymous):

50 is a third of 150, which is 33.33%

OpenStudy (anonymous):

it will be 100/3% to make initial one double.

OpenStudy (anonymous):

let initial =100Rs increased in first year is 25% which is equal to=125Rs in second year it is 20% which is equal to=150Rs now we need to increase 150Rs to make it to 200Rs as initial was 100Rs 100/3 % is the percentage increase.